Initial Setup of MODS1 at Left Direct Gregorian - 19 September 2011 UT

Observer(s): J. Hill (Tucson), R. Pogge, R. Stoll, O. Kuhn (LBTO) plus Fernando, Mauro
Telescope Operator: S. Allanson (LBTO)
Telescope Support: M. Midkiff (LBTO)
Instrument Support: T. Shih (LBTO)
Software Support: T. Leibold (Tucson)

Summary

(These are John's notes from Tucson, which may not reflect all activities happening on the mountain.)

The elevation corrections to the collimation model were installed. They were verified at the LDG focus.

Two important problems with MODS1 guider camera:
1) The CCD image is smeared horizontally. See critical Issue 3652.
2) The guide camera is defocussed compared to the WFS by ~1.5 arcsec.

Beyond those problems, basic functionality at LDG has been checked.

Preparations

Edit into OSU_L.cfg - these were updated later.
   guidecam_ROI   "38 1 479 400"
   guidecam_center_x    213.0
   guidecam_center_y   169.0
   guidecam_hotspot_x   204
   guidecam_hotspot_y   375

John added Andrew's calculated polynomial corrections into SX M2 Collimation model for both LUCI1 and MODS1 (not yet done for IRTC or LBTI)
New from last night
 0.252      2.046     3.000       8.500  -410.100    0.000  Global Offsets
-4.602     -3.652    -1.940      93.600   317.800    0.000  Offset collimation coefficient
 0.08688    0.15323   0.0176      0.725     3.756    0.000  Linear collimation coefficient
-5.633e-4  -0.001876  0.000177    0.03149  -0.028494 0.000  Quadratic collimation coefficient

Old from 20110413
 0.200      1.000    -0.100    -104.000    21.000    0.000  Global Offsets
-3.239     -4.753    -1.940     267.800   427.600    0.000  Offset collimation coefficient
 0.02888    0.1599    0.0176     -1.699    -1.104    0.000  Linear collimation coefficient
-0.0001223 -0.001633  0.000177    0.01709   0.007406 0.000  Quadratic collimation coefficient

Details

Open

opened at sunset - wow!

Checking parameters for the new MODS1 guider camera

02:30 ACQUIRE preset M5 star to IE=-45 CA=+60

02:40 ACQUIRE preset to BS star at EL=72

The lower half of the guide box is cut off (black on the GUI), Torsten confirms that this is because guide box must be contained within guide ROI, but adjusting for that didn't solve the problem.

ROI 38 1 479 450 , and restart GCSL - hangs talking to AzCam WFS Torsten restarts the controller process on mods1-azcamg

Lower half of the guide box is still cut off.

03:10 guideWindowSize=100 (instead of 75) - this worked but not sure why - even numbers?

03:15 center_x = 251, hotspot_x = 242, restart GCSL

These are close to correct values, but the acquisition image graphics are wrong.

Increased box to 140 still didn't show the BS shadow.

03:40 All the guide camera images are smeared horizontally. Rick says this has been an intermittent problem. Cycle power on guide camera controller.

Torsten says the MODS1 and AGW2 cameras use the same DSP and Timing files. We didn't see any hint of this smearing problem with AGw2.

The WFS images are amazingly good running below 100 nm rms. Rick confirms that we have excellent images (0.6 arcsec FWHM) with the MODS1 science detector maybe with some small static aberrations. Maybe the blurring of the guide image (2.3 arcsec FWHM) is displacing the image by an arcsec in the WFS aperture and that makes the aberrations.

03:55 WFS rms is down to 40 nm rms!!!

04:02 Stop WFS loop. Manual focus shows that guide camera is out of focus compared to WFS and MODS1 (by 0.1 mm of M2 Zglobal, Steve moved from Z=2.7 to Z=2.8). See manually focussed guider image 000756. Peak of the in-focus smeared images is 11800, with background of 9800 counts.

04:05 Rebooting the two AzCam computers. Stop and restart GCSL.

04:10 resend the active preset to BS9177 - the guider image is still smeared horizontally.

Changes to OSU_L.cfg
   guidecam_ROI   "38 1 479 450" (maybe should go back to 400?)
   guidecam_center_x    251.0
   guidecam_center_y   169.0
   guidecam_hotspot_x   242
   guidecam_hotspot_y   375
   guideWindowSize   76
Testing the collimation model

Active optics is within 0.6 mm of the collimation model at EL=61. Xact=-0.648 Yact=0.563 Zact=0.075.

04:14 preset to BS star at low elevation.

at EL=35 Xact=-0.505, Yact=0.427, Zact=0.057.

04:23 preset to a star at high elevation

Network connection to the mountain and from lbtdu16 browser to the local twiki is noticeably degraded. ??? Switch to skype audio as polycom won't stay connected.

star at x=32.6 y=-125.2 shows vignetting on WFS. try a different guide star.

Note the entertaining vignetting in left_wfscimage000101. (maybe we are clipping the edge of the WFS aperture here?)

04:43 Clear C00 to get rid of some stray aberrations.

El=76.2 X=-1.329 Y=0.270 Z=0.021 (but maybe not so well collimated as previous, or some off-axis effects)

04:45 Briefly turn on dome lights, but couldn't see much on wfs camera.

Other MODS activities

Exposures on wolf1346, etc

AzCam debugging - swapping parts to confirm where in the chain the problem lies

05:42 Restart GCSL with the fibers swapped between PCI cards in mods1-azcamg.

The problem stayed with the guider camera and did NOT follow the PCI cards.

Rick is writing test images in /OldRepository/20110919_MODS/

05:53 Restore PCI fiber connections to normal.

05:54 preset to verify that function is re-established.

05:59 ACTIVE preset to an on-axis BS star, clear C00 after the first wfs image

06:10 hotspot to 240,377 center to 249,171

selectAGW left OSU_L.cfg

06:13 resend preset

06:27 hotspot to 246,375 center to 255,169

Seeing has degraded to about 2 arcsec. We have an impression that the smearing has gone away, but the bad seeing makes if difficult to be sure. Dome/sky during the reconfiguration show that the smearing is certainly still present. See /OldRepository/20110919_MODS/left_guider000029.fits.

Latest Changes to OSU_L.cfg
   guidecam_ROI   "38 1 479 400"
   guidecam_center_x    255
   guidecam_center_y   169
   guidecam_hotspot_x   246
   guidecam_hotspot_y   375
   guideWindowSize   76

06:35 Move to zenith to reconfigure for LBC

Olga's notes are located in SciPrep20110919

-- JohnHill - 19 Sep 2011
  • LBC Tip Corrections:
    Tip_Corr.jpg

  • LBC Tip Positions:
    Tip_Pos.jpg

  • LBC Tilt Corrections:
    Tilt_Corr.jpg

  • LBC Tilt Positions:
    Tilt_Pos.jpg
Topic attachments
I Attachment Action Size Date Who Comment
Tilt_Corr.jpgjpg Tilt_Corr.jpg manage 48 K 20 Sep 2011 - 20:23 DougMiller LBC Tilt Corrections
Tilt_Pos.jpgjpg Tilt_Pos.jpg manage 39 K 20 Sep 2011 - 20:23 DougMiller LBC Tilt Positions
Tip_Corr.jpgjpg Tip_Corr.jpg manage 52 K 20 Sep 2011 - 20:21 DougMiller LBC Tip Corrections
Tip_Pos.jpgjpg Tip_Pos.jpg manage 51 K 20 Sep 2011 - 20:22 DougMiller LBC Tip Positions
Topic revision: r4 - 20 Sep 2011, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback